OpenStudy (anonymous):

Solve for x. 6x=-14(10+x)

OpenStudy (anonymous):

someone should totally help you! you're awesome and oh so hawt!

OpenStudy (anonymous):

First, open the parentheses: -140 + 14x = 6x Do you want me to continue or you can take it from here?

OpenStudy (anonymous):

Continue! Please & thank you!

OpenStudy (anonymous):

Sorry i made a mistake on my first comment, it's supposed to be: -140 - 14x = 6x

OpenStudy (anonymous):

Now: -14x - 6x = 140 -20x = 140 x = -7

OpenStudy (anonymous):

Oh, I see now! Thank you!

OpenStudy (anonymous):

No Problem.

OpenStudy (anonymous):

<3